# A transversality theorem for semi-algebraic sets with application to signal recovery from the second moment and cryo-EM

### v2

@@ 1 Introduction
A typical cryo-EM experiment produces hundreds of thousands of noisy tomographic projections of the sought molecular structure, each taken from an unknown viewing direction; a detailed mathematical model is introduced in Section 4. The cryo-EM inverse problem is that of estimating the 3-D molecular structure from the noisy projections, while the unknown viewing directions are commonly treated as nuisance variables [9]. Importantly, all existing cryo-EM reconstruction algorithms build upon prior assumptions on the 3-D structure. For example, there is an important class of algorithms, which are based on Bayesian techniques [42, 38]. This paper focuses on the substantial and rich family of semi-algebraic priors.
@@
A semi-algebraic set $\mathcal{M}\subseteq\mathbb{R}^{N}$ is a finite union of sets, which are defined by polynomial equality and inequality constraints. A map $f\colon\mathcal{M}\subset\mathbb{R}^{N}\to\mathcal{N}\subset\mathbb{R}^{M}$ is semi-algebraic if the graph $\Gamma_{f}=\{(x,f(x)\}\subset\mathcal{M}\times\mathcal{N}$ is a semi-algebraic subset of $\mathbb{R}^{N}\times\mathbb{R}^{M}$. The image of a semi-algebraic set under a semi-algebraic map is always semi-algebraic. Any semi-algebraic set $\mathcal{M}$ can be written as a finite union of smooth manifolds, and the dimension of $\mathcal{M}$ is defined to be the maximal dimension of these manifolds. The assumption that a signal (e.g., the 3-D molecular structure) lies in a semi-algebraic set is referred to as a semi-algebraic prior. This work is motivated by three important special cases of semi-algebraic sets.
  • Linear priors: the assumption that the signal of interest lies in some low-dimensional subspace. Linear priors are ubiquitous in signal processing and machine learning and are the root of popular methods, such as Principal Component Analysis (PCA). Linear models were proven highly effective in various stages of the computational pipeline of cryo-EM data processing [48, 52, 49]. Some (but not all) non-linear manifolds are also semi-algebraic sets; a simple example would be a sphere.
@@
    where $z$ resides in a low dimensional (latent) space, the $A_{i}$’s are affine transformations and the $\eta_{i}$’s are semi-algebraic activation functions, such as the rectification function (ReLU) that sets the negative entries of an input to zero. Generative models have demonstrated promising results in various tasks in the cryo-EM computational pipeline and in particular to conformational variability analysis - one of the substantial challenges in the field [53, 20, 37].
@@
In cryo-EM, the method of moments seeks to recover an unknown structure from the higher-order moments of the measurements [9, 34]. For computational and statistical reasons, there is a desire to use moments of minimal order. As proved in [15], when no priors are imposed, the second moment determines a structure up to the action of an ambiguity group $H$, which is a product of certain orthogonal groups. In previous works [16, 15], the authors study sparsity priors that ensure that a structure can be recovered from its second moment. This means that the prior set $\mathcal{M}$ has the property that for any $x\in\mathcal{M}$ the orbit $Hx$ intersects $\mathcal{M}$ only in $x$. In other words, the prior set $\mathcal{M}$ is transverse to the orbits of the ambiguity group $H$.
In this paper, we derive strong recovery guarantees for cryo-EM structures from the second moment, by proving a much more fundamental result. We prove a transversality theorem for semi-algebraic sets in orthogonal or unitary representations $V$ of arbitrary compact Lie groups $H$. In particular, we derive bounds on the dimension of a generic semi-algebraic set $\mathcal{M}$, which ensures that it is transverse to the orbits of the group action; i.e., the $H$ orbit of any point $x\in\mathcal{M}$ intersects $\mathcal{M}$ only at $x$. This implies that if $F$ is a measurement function (such as the second moment) that separates $H$-orbits, then $F$ is one-to-one when restricted to $\mathcal{M}$. This means that the prior knowledge that $x$ lies in $\mathcal{M}$ ensures that $x$ is determined by the measurement $F(x)$.
The main transversality theorems for real orthogonal representations, Theorem 2.2 and Theorem 2.3, are stated in Section 2, and proved in Section 6. Since our primary interest is in real representations, we state and prove the corresponding results for complex unitary representations in Appendix A to avoid cluttering the main body of the text.
@@
The abstract setting of the transversality theorem allows us to derive several corollaries that span a wide range of mathematical models and scientific applications. Our main focus is on recovering signals that are translated by random group actions from their second moment. In Section 3 we introduce the implications of the transversality theorem to this problem and discuss two immediate consequences. First, we derive bounds on the problem of recovering a matrix from its Gram matrix, assuming the matrix lies in a low-dimensional semi-algebraic set. Second, we derive bounds for signal recovery from Fourier magnitudes, under semi-algebraic priors. This problem is called the phase retrieval problem and it plays a key role in a variety of applications in optical imaging, signal processing, and X-ray crystallography. In Section 4 we discuss the implications of the transversality theorems to cryo-EM—our chief motivation. We derive conditions under which a 3-D molecular structure can be recovered from the second moment of the measurements. We show that this implies an improved sample complexity (fewer observations are required for reconstruction) and discuss implications. We also discuss the connection to the multi-reference alignment problem. From a representation theory perspective, cryo-EM and multi-reference alignment can be understood as generalizations of the phase retrieval problem: while in the latter each irreducible representation appears with multiplicity one, the cryo-EM and multi-reference alignment models typically involve multiple copies of each irreducible representation. In Section 5 we apply the transversality theorem to a problem of designing permutation invariant separators in machine learning. This problem does not involve second moments, but rather other group invariants, namely row-wise sorting.
@@
There are many previous works in which transversality results are implicitly proved for specific group actions and for specific types of semi-algebraic priors, such as linear subspaces. For example, the frame phase retrieval results of [5] can be rephrased in the language of this paper as stating that a generic linear subspace of dimension $M$ in $\mathbb{R}^{N}$ with $N>2M-2$ is transverse to the orbits of the group {±$\{\pm 1\}^{N}$ acting on $\mathbb{R}^{N}$ by coordinate sign changes. Because our results apply to any semi-algebraic set, the bounds we obtain from Theorem 2.2 and Theorem 2.3 are sometimes slightly weaker (differing by a constant) from the transversality bounds for specific semi-algebraic sets with simple structures, such as linear subspaces. The relation of our results with the existing phase retrieval literature is discussed further in Section 3. For the cryo-EM case, discussed in Section 4, our bounds are actually much stronger than those previously obtained in the literature for generic orthonormal bases [15]. For the problem of permutations, we also obtain much stronger bounds than previous work [5, 23], (though with a different notion of genericity).
@@ 2 A transversality theorem for semi-algebraic sets
Our notion of a generic semi-algebraic set depends on a choice of a semi-algebraic group $\Theta$ of semi-algebraic automorphisms of $V$. Hereafter, $\theta(\mathcal{M})$ denotes the image of $\mathcal{M}$ under the automorphism $\theta$.
@@
For any semi-algebraic set $\mathcal{M}$ with $\dim\mathcal{M}\leq M$, the transversality result holds for $\theta(\mathcal{M})$, where $\theta$ is a generic element of the semi-algebraic group $\Theta$.
Precisely, we mean that the set of $\theta\in\Theta$ for which the transversality result does not hold for $\theta(\mathcal{M})$ has strictly smaller dimension than $\dim\Theta$.
@@
where $k(H)=\max_{x\in V}\dim Hx$. Note that the orbits of $H$ have dimension [^1] A classical theorem of Chevalley states that any compact Lie group is the set of real points of a complex algebraic group and is, therefore, a real algebraic set. More generally, if $K\subset H$ is a closed (and hence compact) Lie subgroup, then the compact homogeneous space $H/K$ is also a real algebraic set [33, Theorem 3]. It follows that if $V$ is a representation of a compact Lie group $H$, then the orbit $Hx$ of any $x\in V$ is also a real algebraic set since the orbit is isomorphic to the compact homogeneous space $H/H_{x}$ where $H_{x}=\{h\in H|h\cdot x=x\}$. In particular, when we talk about the dimension of the orbits, we are considering them as real algebraic subsets of $V$. at most equal to $\dim V$ so $k(H)\leq\dim V$ and hence $K\geq 0$. Clearly, the problem gets easier for smaller $M$ (the semi-algebraic set is of low-dimension) and larger $K$ (the effective dimension of the representation is larger); we prove that the gap between $K$ and $M$ can be small. The theorems are proved in Section 6.
@@
    If $K>M$, then for a generic $x$ in a $\mathrm{GL}(V)$ or $\mathrm{Aff}$-generic semi-algebraic subset $\mathcal{M}$ of dimension $M$, if $h\cdot x\in M$ for some $h\in H$, then $h\cdot x=\pm x$ (for $\mathrm{GL}(V)$) or $h\cdot x=x$ (for $\mathrm{Aff}(V)$).
@@
    If $K>2M$, then for all $x$ in a $\mathrm{GL}(V)$ or $\mathrm{Aff}$-generic semi-algebraic subset $\mathcal{M}$ of dimension $M$, if $h\cdot x\in\mathcal{M}$ for some $h\in H$, then $h\cdot x=\pm x$ (for $\mathrm{GL}(V)$) or $h\cdot x=x$ (for $\mathrm{Aff}(V)$).
An illustration of this theorem when $V=\mathbb{R}^{2}$ is provided in Figure 1. Panel (a) depicts an $M=0$ dimensional semialgebraic set $\mathcal{M}$, which we think of as the generic affine image of an axis-aligned finite grid. We consider the orbit of points in this set (like the point denoted by the green arrow), under the action of $\{-1,1\}^{2}$ by elementwise multiplication. Since $2=K>0=2M$, the theorem guarantees that the orbits will not intersect the grid. In panel (b) we consider the same group action, now on a one-dimensional semi-algebraic set, a union of a line and a circle. In this case, $K=2$, which is larger than $M=1$ but not larger than $2M$. Therefore, the theorem guarantees that orbits of generic points (like the point marked by a green arrow) will not intersect the set, but it is possible that orbits of a measure zero set of points, like the point denoted by a red arrow, will intersect the set. Finally, in panel (c), we consider an $M=1$ dimensional semi-algebraic set, and the orbits of $S^{1}$, a one-dimensional group. In this case, we have $K=1=M$, so the conditions of the theorem are violated. Indeed, the orbits of all non-zero points have a non-trivial intersection.
Figure 1: Panels (a)-(c) show three different semi-algebraic sets, colored blue: a zero-dimensional grid in (a), a union of a line and circle in (b), and a union of two lines in (c). Each panel selects points in the semi-algebraic set (denoted by red or green arrows) and shows their orbits under the zero-dimensional group $\mathbb{Z}_{2}^{2}$ (for (a) and (b)) or the one-dimensional group $S^{1}$ in (c). In (a) the orbits do not intersect the semi-algebraic set, in (b) most orbits do not intersect, but the orbits of a measure zero set of points (like the point denoted by the red arrow) will intersect the set. Intersections in (c) are inevitable at all points except for zero. The different behavior in these three cases is determined (generically) by the dimension of the group orbit, the dimension of the semialgebraic set, and the ambient dimension, see Theorem 2.2.
@@
If in addition the orbits of $H$ are connected (for example if $H$ is connected) then we obtain the improved bounds of $K>M+1$ and $K>2M+1$ respectively.
@@
The results stated above depend on the dimension of the orbit $Hx$. If $H$ is a finite group, then the dimension of the orbit is always zero. If $H$ is positive-dimensional, then the orbits of $H$ can have dimension strictly smaller than $\dim H$. For example, if $H=\operatorname{GL}(V)$ acting by linear automorphisms on an $N$-dimensional (real) vector space $V$, then $\dim H=N^{2}$ but non-zero vectors all lie in a single orbit of dimension equal to $N=\dim V$, and the origin is its own orbit. If $H=\mathrm{O}(V)$, then $\dim H=\binom{N}{2}$, and the orbits of non-zero vectors are spheres and thus have dimension $N-1$.
@@
The bound $K>M$ in Theorem 2.2 is sharp, since if $H$ is finite, then $K=\dim V$. And if we take $\mathcal{M}=V$, then clearly the conclusion of the theorem cannot hold for any translate of $V$ since all translates are equal to $V$.
@@
The following examples show that the bounds $K>M+2$ and $K>M+1$ in Theorem 2.3 are also sharp. Let =$\mathcal{M}=\{(1,0),(-1,0),(0,1),(0,-1)\}\subset\mathbb{R}^{2}$. If we take $H=\mathbb{Z}_{4}$, where the generator of $\mathbb{Z}_{4}$ acts by $(x,y)\mapsto(y,-x)$, then $K=2$ and $\dim\mathcal{M}=0$, but any rotation of $\mathcal{M}$ produces four points which lie in a single $H$ orbit. Similarly, if we take $H=\operatorname{SO}(2)$, then the orbits of $H$ are connected and one-dimensional so $K=1$, but again any rotation $\mathcal{M}$ consists of points lying in the same $H=\operatorname{SO}(2)$ orbit.

### v3

@@ 1 Introduction
A typical cryo-EM experiment produces hundreds of thousands of noisy tomographic projections of the sought molecular structure, each taken from an unknown viewing direction; a detailed mathematical model is introduced in Section 4. The cryo-EM inverse problem involves estimating the 3-D molecular structure from noisy projections, while the unknown viewing directions are commonly treated as nuisance variables [10]. Importantly, all existing cryo-EM reconstruction algorithms build upon prior assumptions on the 3-D structure. For example, there is an important class of algorithms, which are based on Bayesian techniques [45, 40]. This paper focuses on the substantial and rich family of semi-algebraic priors.
@@
A semi-algebraic set $\mathcal{M}\subseteq\mathbb{R}^{\color[rgb]{0,0,0}S}$ is a finite union of sets, which are defined by polynomial equality and inequality constraints. If $\mathcal{M}\subset\mathbb{R}^{S},\mathcal{N}\subset\mathbb{R}^{T}$ are semi-algebraic sets, we will say that a map $f\colon\mathcal{M}\to\mathcal{N}$ is semi-algebraic if the graph $\Gamma_{f}=\{(x,f(x)\}\subset\mathcal{M}\times\mathcal{N}$ is a semi-algebraic subset of $\mathbb{R}^{S}\times\mathbb{R}^{T}$. The image of a semi-algebraic set under a semi-algebraic map is always semi-algebraic. Any semi-algebraic set $\mathcal{M}$ can be written as a finite union of smooth manifolds, and the dimension of $\mathcal{M}$ is defined to be the maximal dimension of these manifolds. The assumption that a signal (e.g., the 3-D molecular structure) lies in a semi-algebraic set is referred to as a semi-algebraic prior. This work is motivated by three important special cases of semi-algebraic sets.
  • Linear priors: the assumption that the signal of interest lies in some low-dimensional subspace. Linear priors are ubiquitous in signal processing and machine learning, and they are the foundation of popular methods, such as Principal Component Analysis (PCA). Linear models were proven highly effective in various stages of the computational pipeline of cryo-EM data processing [53, 57, 54]. Some (but not all) non-linear manifolds are also semi-algebraic sets; a simple example would be a sphere.
@@
    where $z$ resides in a low-dimensional (latent) space, the $A_{i}$’s are affine transformations and the $\eta_{i}$’s are semi-algebraic activation functions, such as the rectification function (ReLU) that sets the negative entries of an input to zero. Generative models have demonstrated promising results in various tasks within the cryo-EM computational pipeline, particularly in conformational variability analysis - one of the substantial challenges in the field [58, 21, 39].
@@
In cryo-EM, the method of moments seeks to recover an unknown structure from the higher-order moments of the measurements [10, 35]. For computational and statistical reasons, there is a desire to use moments of minimal order. As proved in [16], when no priors are imposed, the second moment determines a structure up to the action of an ambiguity group $H$, which is a product of certain orthogonal groups. In previous works [17, 16], the authors study sparsity priors that ensure that a structure can be recovered from its second moment. This means that the prior set $\mathcal{M}$ has the property that for any $x\in\mathcal{M}$ the orbit $Hx$ (that is, the orbit under the group of the product of orthogonal matrices) intersects $\mathcal{M}$ only in $x$. In other words, the prior set $\mathcal{M}$ is transverse to the orbits of the ambiguity group $H$.
In this paper, we derive strong recovery guarantees for cryo-EM structures from the second moment by proving a much more fundamental result. We prove a transversality theorem for semi-algebraic sets in orthogonal or unitary representations $V$ of arbitrary compact Lie groups $H$. In particular, we derive bounds on the dimension of a generic translate of a semi-algebraic set $\mathcal{M}$, which ensures that it is transverse to the orbits of the group action; i.e., the $H$ orbit of any point $x\in\mathcal{M}$ intersects $\mathcal{M}$ only at $x$. This implies that if $F$ is a measurement function (such as the second moment) that separates $H$-orbits, then $F$ is one-to-one when restricted to $\mathcal{M}$. This means that the prior knowledge that $x$ lies in $\mathcal{M}$ ensures that $x$ is determined by the measurement $F(x)$. The main transversality theorems for real orthogonal representations, Theorem 2.2 and Theorem 2.3, are stated in Section 2, and proved in Section 6. Since our primary interest is in real representations, we state and prove the corresponding results for complex unitary representations in Appendix A to avoid cluttering the main body of the text.
Remark 1.1.
We note that our use of the term transversality is not the usual differential-geometric notion of transversal intersection of manifolds. Instead, we were inspired by S. Kleiman’s celebrated theorem in his paper The transversality of a general translate [36]. There, he proves that if $V,W$ are two subvarieties of a homogeneous space for some algebraic group $H$, then the general translate of $V$ by an element of $H$ intersects $W$ in the expected dimension. While the setup of our transversality results is quite different from Kleiman’s, our main results are in the spirit of Kleiman’s classic theorem.
@@
The abstract setting of the transversality theorem allows us to derive several corollaries that span a wide range of mathematical models and scientific applications. Our main focus is on recovering signals that are translated by random group actions from their second moment. In Section 3, we introduce the implications of the transversality theorem to this problem and discuss two immediate consequences. First, we derive bounds on the problem of recovering a matrix from its Gram matrix, assuming the matrix lies in a low-dimensional semi-algebraic set. Second, we derive bounds for signal recovery from Fourier magnitudes, under semi-algebraic priors. This problem is called the phase retrieval problem, and it plays a key role in a variety of applications in optical imaging, signal processing, and X-ray crystallography. In Section 4 we discuss the implications of the transversality theorems to cryo-EM—our chief motivation. We derive conditions under which a 3-D molecular structure can be recovered from the second moment of the measurements. We show that this implies an improved sample complexity (fewer observations are required for reconstruction) and discuss implications. We also discuss the connection to the multi-reference alignment problem. From a representation theory perspective, cryo-EM and multi-reference alignment can be understood as generalizations of the phase retrieval problem: While in the latter each irreducible representation appears with multiplicity one, the cryo-EM and multi-reference alignment models typically involve multiple copies of each irreducible representation. In Section 5 we apply the transversality theorem to a problem of designing permutation invariant separators in machine learning. This problem does not involve second moments, but rather other group invariants, namely row-wise sorting.
@@
There are many previous works in which transversality results are implicitly proved for specific group actions and for specific types of semi-algebraic priors, such as linear subspaces. For example, the frame phase retrieval results of [5] can be rephrased in the language of this paper as stating that a generic linear subspace of dimension $M$ in $\mathbb{R}^{N}$ with N>2M2N>2M-2 is transverse to the orbits of the group $\{\pm 1\}^{N}$ acting on $\mathbb{R}^{N}$ by coordinate sign changes. Because our results apply to any semi-algebraic set, the bounds we obtain from Theorem 2.2 and Theorem 2.3 are sometimes slightly weaker (differing by a constant) from the transversality bounds for specific semi-algebraic sets with simple structures, such as linear subspaces. The relation of our results with the existing phase retrieval literature is discussed further in Section 3. For the cryo-EM case, discussed in Section 4, our bounds are actually much stronger than those previously obtained in the literature for generic orthonormal bases [16]. For the problem of permutations, we also obtain much stronger bounds than previous work [5, 24], (though with a different notion of genericity).
@@ 2 A transversality theorem for semi-algebraic sets
Our notion of a generic semi-algebraic set depends on a choice of a semi-algebraic group $\Theta$ of semi-algebraic automorphisms of $V$. Hereafter, $\theta\cdot\mathcal{M}$ denotes the image of $\mathcal{M}$ under the automorphism $\theta$.
@@
For any semi-algebraic set $\mathcal{M}$ with $\dim\mathcal{M}\leq M$, the transversality result holds for $\theta\cdot\mathcal{M}$, where $\theta$ is a generic element of the semi-algebraic group $\Theta$.
Precisely, we mean that the set of $\theta\in\Theta$ for which the transversality result does not hold for $\theta\cdot\mathcal{M}$ has strictly smaller dimension than $\dim\Theta$.
@@
where $k(H)=\max_{x\in V}\dim Hx$. Note that the orbits of $H$ have dimension [^1] A classical theorem of Chevalley states that any compact Lie group is the set of real points of a complex algebraic group and is, therefore, a real algebraic set. More generally, if $K\subset H$ is a closed (and hence compact) Lie subgroup, then the compact homogeneous space $H/K$ is also a real algebraic set [34, Theorem 3]. It follows that if $V$ is a representation of a compact Lie group $H$, then the orbit $Hx$ of any $x\in V$ is also a real algebraic set, since the orbit is isomorphic to the compact homogeneous space $H/H_{x}$ where $H_{x}=\{h\in H|h\cdot x=x\}$. In particular, when we talk about the dimension of the orbits, we are considering them as real algebraic subsets of $V$. at most equal to $\dim V$ so $k(H)\leq\dim V$ and hence $K\geq 0$. Clearly, the problem gets easier for smaller $M$ (the semi-algebraic set is of low dimension) and larger $K$ (the effective dimension of the representation is larger); we prove that the gap between $K$ and $M$ can be small. The theorems are proved in Section 6.
@@
    If $K>M$, then for a generic $x$ in a $\mathrm{GL}(V)$ or $\mathrm{Aff}(V)$-generic semi-algebraic subset $\mathcal{M}$ of dimension $M$, if $h\cdot x\in{\color[rgb]{0,0,0}\mathcal{M}}$ for some $h\in H$, then $h\cdot x=\pm x$ (for $\mathrm{GL}(V)$) or $h\cdot x=x$ (for $\mathrm{Aff}(V)$).
@@
    If $K>2M$, then for all $x$ in a $\mathrm{GL}(V)$ or $\mathrm{Aff}(V)$-generic semi-algebraic subset $\mathcal{M}$ of dimension $M$, if $h\cdot x\in\mathcal{M}$ for some $h\in H$, then $h\cdot x=\pm x$ (for $\mathrm{GL}(V)$) or $h\cdot x=x$ (for $\mathrm{Aff}(V)$).
An illustration of this theorem when $V=\mathbb{R}^{2}$ is provided in Figure 1. Panel (a) depicts an $M=0$ dimensional semialgebraic set $\mathcal{M}$, which we think of as the generic affine image of an axis-aligned finite grid. We consider the orbit of points in this set (like the point denoted by the green arrow), under the action of $\{-1,1\}^{2}$ by elementwise multiplication. Since $2=K>0=2M$, the theorem guarantees that the orbits will only have a trivial intersection with the grid (namely, only the point itself is both in its orbit and in the set). In panel (b), we consider the same group action, now on a one-dimensional semi-algebraic set, a union of a ray and a circle. In this case, $K=2$, which is larger than $M=1$ but not larger than $2M$. Therefore, the theorem guarantees that orbits of generic points (like the point marked by a green arrow) will not intersect the set non-trivially, but it is possible that orbits of a measure-zero set of points, like the point denoted by a red arrow, will have a non-trivial intersection with the set. Finally, in panel (c), we consider an $M=1$ dimensional semi-algebraic set, and the orbits of $S^{1}$, a one-dimensional group. In this case, we have $K=1=M$, so the conditions of the theorem are violated. Indeed, the orbits of all non-zero points have a non-trivial intersection.
Figure 1: Panels (a)-(c) show three different semi-algebraic sets, colored blue: a zero-dimensional grid in (a), a union of a ray and a circle in (b), and a union of two lines in (c). Each panel selects points in the semi-algebraic set (denoted by red or green arrows) and shows their orbits under the zero-dimensional group $\mathbb{Z}_{2}^{2}$ (for (a) and (b)) or the one-dimensional group $S^{1}$ in (c). In (a), the orbits do not have a non-trivial intersection with the semi-algebraic set, in (b) most orbits have only a trivial intersection, but the orbits of a measure zero set of points (like the point denoted by the red arrow) will have a non-trivial intersection. Non-trivial intersections in (c) are inevitable at all points except for zero. The different behavior in these three cases is determined (generically) by the dimension of the group orbit, the dimension of the semialgebraic set, and the ambient dimension, see Theorem 2.2.
@@
If, in addition, the orbits of $H$ are connected (for example, if $H$ is connected), then we obtain the improved bounds of $K>M+1$ and $K>2M+1$ respectively.
@@
The results stated above depend on the dimension of the orbit $Hx$. If $H$ is a finite group, then the dimension of the orbit is always zero. If $H$ is positive-dimensional, then the orbits of $H$ can have dimension strictly smaller than $\dim H$. For example, if $H=\operatorname{GL}(V)$ acting by linear automorphisms on an $N$-dimensional (real) vector space $V$, then $\dim H=N^{2}$, but non-zero vectors all lie in a single orbit of dimension equal to $N=\dim V$, and the origin is its own orbit. If $H=\mathrm{O}(V)$, then $\dim H=\binom{N}{2}$, and the orbits of non-zero vectors are spheres and thus have dimension N1N-1.
@@
The bound $K>M$ in Theorem 2.2 is sharp, since if $H$ is finite, then $K=\dim V$. And if we take $\mathcal{M}=V$, then clearly the conclusion of the theorem cannot hold for any translate of $V$ since all translates are equal to $V$. Here is an example to show that the bound $K>2M$ in the second part of the theorem is sharp. Consider the group $H=\{\pm{1}\}^{2}$ acting by coordinate-wise sign change. If $\mathcal{M}$ is the parabola defined by the equation $y-x^{2}=0$, then a $\operatorname{GL}(V)$-translate of $\mathcal{M}$ is a parabola with equation $ax+by-(cx+dy)^{2}=0$, where $ad-bc\neq 0$. There is a set of $(a,b,c,d)$ of positive measure in $\mathrm{GL}(V)$ where this parabola contains real points of the form (x,y),(x,y)(x,y),(-x,y) with $xy\neq 0$. For example, taking $(a,b,c,d)=(1,2,2,4)$, the points $(\pm\frac{1}{8},\frac{1}{16})$ both lie on the parabola. Likewise, a translate of $\mathcal{M}$ by an element of $\mathrm{Aff}(V)$ has equation $ax+by+A-(cx+dy+B)^{2}=0$, where $A,B$ are arbitrary scalars. Once again there is a set of positive measure in $\mathrm{Aff}(V)$ where the translated parabola has real points of the forms (x,y),(x,y)(x,y),(-x,y) with $xy\neq 0$.
@@
The following examples show that the bounds $K>M+2$ (resp. $2M+2$) and $K>M+1$ (resp. $2M+1$) in Theorem 2.3 are also sharp. Let $\mathcal{M}=\{(1,0),(-1,0),(0,1),(0,-1)\}\subset\mathbb{R}^{2}$. If we take $H=\mathbb{Z}_{4}$, where the generator of $\mathbb{Z}_{4}$ acts by $(x,y)\mapsto(y,-x)$, then $K=2$ and $\dim\mathcal{M}=0$, but any rotation of $\mathcal{M}$ produces four points which lie in a single $H$ orbit. Similarly, if we take $H=\operatorname{SO}(2)$, then the orbits of $H$ are connected and one-dimensional so $K=1$, but again any rotation $\mathcal{M}$ consists of points lying in the same $H=\operatorname{SO}(2)$ orbit. Note that for a prior $\mathcal{M}$ of dimension zero transversality holds at a generic point of $\mathcal{M}$ if and only if it holds at all points.
The implication of the genericity assumption.
The main results of this paper rely on the assumption that the signal lies in a generic translate of a low-dimensional semi-algebraic set. While this implies that the results hold for almost every such translate, they do not necessarily apply to a fixed given semi-algebraic set. To illustrate this, consider the three primary examples introduced in Section 1: linear priors, sparsity, and deep generative models.
In the case of deep generative models, the genericity assumption is typically not a major limitation, as it can often be satisfied by appending a generic (e.g., random) layer to a fixed neural network. For the sparsity assumption, if the sparse basis is learned from data—as in dictionary learning [26]—it is reasonable to expect the model to be generic. However, when sparsity is defined with respect to a fixed, application-driven basis, our results no longer apply. A prominent example is X-ray crystallography, where signals are sparse in the standard basis. Such cases require alternative analytical tools [27, 14]; see further discussion in Section 3.3.2.
The same caveat applies to linear priors. If the signal resides in a low-dimensional learned subspace, such as one obtained via PCA, the genericity assumption is likely to hold. In contrast, if the subspace is fixed, as in coherent diffraction imaging [9], our theoretical results do not apply.
Context of the transversality theorem.
As noted above, previous results in phase retrieval, such as frame phase retrieval, can be viewed as transversality theorems for the intersections of linear subspaces with group orbits. To the best of our knowledge, general results like Theorems 2.2, 2.3 have not previously appeared in the math literature. However, our work is inspired by classical moving lemmas, which were proved in intersection theory [43]. In algebraic geometry,a moving lemma states that if $\alpha$ and $\beta$ are projective cycles [^2] A cycle of dimension $k$ is a finite formal sum of $k$-dimensional subvarieties. in the projective space $\mathbb{P}^{n}$ of dimensions $k$ and $\ell$, respectively, then the cycle $\alpha$ can be moved in its rational equivalence class to a cycle $\alpha^{\prime}$ so that $\alpha^{\prime}$ and $\beta$ intersect in a cycle of the expected dimension $k+\ell-n$ (or is empty if $k+\ell-n$ is negative). In the context of group actions, Kleiman’s transversality theorem [36] states that if $V$ and $W$ subvarieties of the orbit $X$ of an algebraic group $H$, then for a general element $\sigma\in H$ the translate $\sigma\cdot V$ intersects $W$ in dimension $\dim V+\dim W-\dim X$, or is empty if this number is negative.
Here, we consider a different problem—that of identifying translates of a semi-algebraic set with a transversality property with respect to the orbits of a compact group $H$. However, the idea of finding a translate of a given set with desired properties is inspired by these older theorems. As is in the case with Kleiman’s theorem, we seek to translate by an automorphism of the ambient space, which for us is the representation $V$ of $H$. The specific results depend on the particular choice of automorphisms, and we have focused on the groups $\mathrm{Aff}(V)$, $\mathrm{GL}(V)$, and $\operatorname{SO}(V)$ because they naturally occur in applications of interest.